OpenStudy (anonymous):

what is meant by rationalization and how can i rationalize the denominator explain with examples

OpenStudy (anonymous):

okay rationalization is for making denominator real \[\frac{2}{\sqrt{5}}\] \[\frac{2}{\sqrt5} \times \frac{\sqrt5}{\sqrt5}\] \[\frac{2 \sqrt5}{5}\]

OpenStudy (angela210793):

example: 2/sqr6 multiply up and down with sqr6-->2sqr6/sqr6*sqr6=2sqr6/6=sqr6/3

OpenStudy (anonymous):

thnk u